# Large Language Models for Software Engineering: A Systematic Literature Review

### v5

@@ 2.3.2. Study Quality Assessment
A well-crafted quality assessment can help to prevent biases introduced by low-quality studies and can indicate to readers where caution about conclusions should be drawn (Yang et al., 2021). We formulated ten Quality Assessment Criteria (QAC), as shown in Table 4. These aim to assess the relevance, clarity, validity, and significance of included papers. We used a scoring system of -1, 0, 1 (irrelevant/unmet, partially relevant/met, relevant/fully met). The first three questions were designed for the remaining 382 papers in the fifth stage. If QAC1, QAC2, or QAC3 received a score of -1, there is no need to proceed with QAC4-QAC10, and the paper can be excluded directly. QAC4-QAC10 involved assessing the content of the papers using a scoring system of 0, 1, 2, 3 (poor, fair, good, excellent). Finally, we calculated the total score of QAC4-QAC10 for each paper. For published papers, the maximum score for QAC4-QAC10 should be 21 (3 $\times$ 7). We retained papers with a score of 16.8 (21 $\times$ 0.8) or above. For unpublished papers on arXiv, the score for QAC4 is always 0, and the maximum score for QAC5-QAC10 should be 18 (3 $\times$ 6). We retained papers with a score of 14.4 (18 $\times$ 0.8) or above. After this quality assessment, we obtained a final set of 218 papers.
@@ 2.4. Snowballing Search
Before conducting snowballing, a set of initial papers needs to be prepared. In this study, the initial paper list consists of the remaining 218 papers after the quality assessment. We performed forward and backward snowballing, which resulted in the collection of 2,034 and 6,829 papers, respectively. After initial deduplication, we were left with 3,350 papers. We then conducted the full study selection process on these 3,350 papers, including deduplicating them with the 218 papers from performing snowballing on the initial list. As a result, we obtained an additional 11 papers.
@@ 2.5. Data Extraction and Analysis
Figure 2. Overview of the selected 229 papers’ distribution.
We finally obtained 229 relevant research papers after searching and snowballing. Fig. 2 presents an overview of the distribution of the included papers. As shown in Fig. 2 (a), 38% of papers are published in peer-reviewed venues. ICSE is the most common of these venues, with a contribution of 9% of the total. Other venues with noteworthy contributions include TSE, ICSME, and SANER, contributing 5%, 3%, and 3% respectively. Meanwhile, the remaining 62% of papers are published on arXiv, an open-access platform that serves as a repository for scholarly articles. This finding is not surprising since much new LLM4SE research is rapidly emerging and thus many works are just completed and are likely in the peer review process. Despite the non-peer-reviewed nature of these papers, we have performed a rigorous quality assessment process on all collected papers, to ensure the quality of validity of our findings. This approach allows us to include all high-quality and relevant publications while maintaining high research standards.
Fig. 2 (b) shows the temporal distribution of the included papers. The number of publications has seen a rapidly growing trend since 2020. In 2020 and 2021, there are only 7 and 11 relevant papers, respectively. However, by 2022, the number of papers increases dramatically to 51. What’s surprising is that, in the first half of 2023 alone, the number of published papers has already reached 160. This rapid growth trend demonstrates that there is a growing research interest in the domain of LLM4SE.
In order to visualize the main content of our collection of papers, we generated a word cloud based on the abstracts of 229 papers as shown in Fig. 3. The most frequently occurring words include “code”, “LLM”, “task”, “generation”, “performance”, and “program”, clearly indicating the main themes explored in these papers. The term “code” emphasizes the core elements of software engineering, while “LLM” denotes the use of large language models in a variety of tasks. The terms “generation” and “task” emphasize the use of the LLM for automatic code generation and other SE tasks. In addition, “performance” reflects the evaluation and assessment of the effectiveness of LLM in SE applications. The word cloud provides further visual evidence that the literature we have collected is closely related to our research topic, which is to investigate the application of LLM in SE tasks.
@@
We then conducted data extraction during the full-text review. This extraction phase collected all relevant data that would facilitate a comprehensive and insightful response to the RQs outlined in Section 2.1. As depicted in Table 5, we extracted data including the classification of SE tasks, their corresponding activities, as well as the category, characteristics, and applicability of the LLMs. With this collected data, we systematically analyzed the relevant aspects of LLM application in the SE domain.
@@ 3.1. Large Language Models (LLMs)
Pre-trained language models (PLMs) have demonstrated impressive capabilities in solving various NLP tasks (Kojima et al., 2022; Shanahan, 2022; Wei et al., 2022b; Zhao et al., 2023c). Researchers have observed that scaling up the model sizes significantly enhances their capacity, leading to remarkable performance improvements when the parameter scale surpasses a certain threshold (Shanahan, 2022; Hoffmann et al., 2022; Taylor et al., 2022). The term “Large Language Model” (LLM) was introduced to distinguish language models based on their parameter size, specifically referring to large-sized PLMs (Zhao et al., 2023c). However, we note that the literature lacks a formal consensus on the minimum parameter scale for LLMs, as the model’s capacity is intertwined with both data size and total compute (Wang et al., 2023a). In this paper, we adopt the LLM scope division and taxonomy introduced by Pan et al.(Pan et al., 2023b) and categorize the mainstream LLMs investigated in this study into three groups according to their architectures: encoder-only, encoder-decoder, and decoder-only LLMs. This taxonomy and relevant models are shown in Fig. 4.
@@
Encoder-only LLMs. Encoder-only LLMs are a type of neural network architecture that utilizes only the encoder component of the model (Devlin et al., 2018). The encoder’s function is to process and encode the input sentence into a hidden representation, capturing the relationships between words and the overall context of the sentence. Notable instances of encoder-only LLMs include BERT (Devlin et al., 2018) and its variants (Feng et al., 2020; Guo et al., 2020; Liu et al., 2019; Lan et al., 2019). As an example, BERT’s structure, based on the Transformer’s encoder architecture, has been referenced in 41 of the papers in this study. Its distinctive bidirectional attention mechanism simultaneously considers the left and right context of each word during training. In the SE domain, other prominent models like CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2020), RoBERTa (Liu et al., 2019), and ALBERT (Lan et al., 2019) have been widely employed. Specialized models such as BERTOverflow (Tabassum et al., 2020) and CodeRetriever (Li et al., 2022b) have been specifically developed for SE applications. These models’ innovations differ from BERT by leveraging the program structure, introducing new pre-training tasks, or engaging new modalities, thereby improving the architecture’s application to code-related tasks. For example, CodeBERT integrates a token prediction scheme to comprehend code by predicting subsequent tokens, enhancing its understanding of programming languages for tasks like code completion and bug detection (Feng et al., 2020). GraphCodeBERT introduces edge-type prediction, recognizing relationships between code elements as a graph. This enables GraphCoderBERT to leverage code structure, improving its effectiveness in tasks like code summarization and program analysis (Guo et al., 2020). These models have shown efficacy in tasks requiring a nuanced understanding of the entire sentence or code snippet. Examples include code review, bug report understanding, and named entity recognition pertaining to code entities (Pudari and Ernst, 2023; Sghaier and Sahraoui, 2023; Yang et al., 2022c; Arakelyan et al., 2023; Li et al., 2023c; Mukherjee and Hellendoorn, 2023).
@@
Encoder-decoder LLMs. Encoder-decoder LLMs incorporate both encoder and decoder modules (Vaswani et al., 2017). The encoder ingests the input sentence and encodes it into a hidden space, effectively capturing the underlying structure and semantics. This hidden representation serves as an intermediary language, bridging the gap between diverse input and output formats. Conversely, the decoder utilizes this hidden space to generate the target output text, translating the abstract representation into concrete and contextually relevant expressions. Models such as PLBART (Ahmad et al., 2021), T5 (Raffel et al., 2020), and CodeT5 (Wang et al., 2021a) embody this architecture. Further advancements are evident in CodeT5+ (Wang et al., 2023c), while AlphaCode (Li et al., 2022a) and CoTexT (Phan et al., 2021) showcase the architecture’s adaptability to various SE tasks. The encoder-decoder design offers flexible training strategies and is proficient in handling multifaceted tasks such as summarization, translation, and question-answering. Within the field of SE, this ability has been successfully applied to tasks like code summarization (Al-Kaswan et al., 2023; Gu et al., 2022; Mastropaolo et al., 2021b). The encoder module’s capacity to understand and represent both the structure and semantics of code is pivotal, allowing the decoder to translate this comprehension into concise, human-readable summaries.
@@
Decoder-only LLMs. Decoder-only LLMs exclusively utilize the decoder module to generate the target output text, following a distinct training paradigm that emphasizes sequential prediction (Radford et al., 2018). Unlike the encoder-decoder architecture, where the encoder processes input text, the decoder-only architecture begins with an initial state and predicts subsequent tokens, gradually building the output text. This approach relies heavily on the model’s ability to understand and anticipate language structure, syntax, and context. GPT-series models, such as GPT-1 (Radford et al., 2018), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), GPT-3.5 (OpenAI, 2022b), GPT-4 (OpenAI, 2023b), as well as their notable derivative, ChatGPT (OpenAI, 2022a) [^2] ChatGPT is a conversational agent built upon the GPT architecture, with GPT-3.5 and GPT-4 being specific versions of the architecture, each representing successive advancements., represent their major implementations. More specialized versions like CodeGPT (Lu et al., 2021), InstructGPT (Ouyang et al., 2022), Codex (Chen et al., 2021b), Copilot (GitHub, 2023) [^3] Copilot is an application built upon LLMs tailored for coding tasks. For convenience, all subsequent references in this paper to LLMs and their applications, such as ChatGPT and Copilot, will collectively be referred to as LLMs., and others have been fine-tuned for specific tasks in SE. Open-source models like GPT-J (Wang and Komatsuzaki, 2021), GPT-Neo (Black et al., 2021), GPT-NeoX (Black et al., 2022), LLaMA (Touvron et al., 2023a), and Vicuna (Chiang et al., 2023) also follow this architecture. These models can generally perform downstream tasks from a few examples or simple instructions without adding prediction heads or fine-tuning, making them valuable tools in SE research. The year 2022 marked a surge in the development of decoder-only LLMs, a trend that gained further momentum in 2023, notably with the launch of commercial products by leading Internet companies. For example, Google launched Bard (Google, 2023), Meta introduced LLaMA (Touvron et al., 2023a) and Llama 2 (Touvron et al., 2023b), Microsoft unveiled Bing Chat (Microsoft, 2023), etc. Contrary to LLMs such as GPT-4 and its derivative application, ChatGPT, released by OpenAI, which were promptly integrated into SE tasks, these new additions have not yet found widespread application within the SE field. Their potential remains largely unexplored, with opportunities for further assessment and utilization in specific tasks and challenges. The continued advancement of these models emphasizes the active exploration and innovation within decoder-only architectures.
@@ 3.2. Trend Analysis
As shown in Fig. 5, in the span from 2020 to 2023, the architecture of LLMs has witnessed notable shifts in preference and application within SE tasks. The specific choices between decoder-only, encoder-decoder, and encoder-only structures have shaped the direction of research and solutions in the SE domain (Wong et al., 2023). This analysis explores trends in the adoption of these architectures over the years, reflecting the evolving dynamics of LLM for SE tasks.
@@
Evolution of LLM architectures in 2021. The year 2020 saw research papers predominantly concentrating on encoder-only LLMs for SE tasks, evidenced by a total of eight papers. Decoder-only LLMs or encoder-decoder LLMs were not featured in that year’s research. A marked change occurred in 2021. Out of 15 papers in 2021, five were dedicated to decoder-only LLMs, constituting 33.33% of the research. Additionally, three papers, or 20%, focused on encoder-decoder LLMs. Encoder-only LLMs witnessed a slight decline, representing 46.67% of the field with seven papers. This rapid transition can be linked to the generative capability of decoder-only LLMs. Researchers (Laskar et al., 2023; Sadik et al., 2023; Sridhara et al., 2023) found that these models, e.g., GPT series, requiring minimal fine-tuning, could produce not only syntactically correct but also functionally relevant code snippets. Their proficiency in grasping the context of code quickly made them a preferred choice.
@@
Diversity of LLM architectures in 2022. 2022 experienced a significant increase in diversity, with more varied LLM architectures finding representation. Out of a total of 112 papers, 47 were centered around decoder-only LLMs, comprising 41.96% of the studies. Encoder-decoder LLMs made their presence known in 16 papers, accounting for 14.29%. Meanwhile, encoder-only LLMs led the field slightly with 49 papers, capturing 43.75% of the research interest. This diverse distribution suggests an exploration phase where researchers were actively assessing and leveraging different architectures to suit varied needs and challenges. The near-equal interest across different architectures underscores the field’s richness, indicating that no single approach had become the definitive choice.
@@
Dominance of the decoder-only architecture in 2023. 2023 signaled a strong shift towards decoder-only LLMs. An impressive 258 instances of utilizing decoder-only LLMs were recorded across 138 unique papers, reflecting that a single paper might employ multiple such models. These papers focusing on decoder-only LLMs constituted a significant 73.09% of the total research this year. In comparison, encoder-decoder LLMs were the subject of 40 papers, contributing 11.33%, while encoder-only LLMs appeared to stabilize, with 55 papers, representing 15.58% of the 2023 research landscape. This trend signifies a shift in focus and resources toward exploring and harnessing the decoder-only architecture as the primary approach in many current and future LLM4SE research and applications.
@@
Criteria for LLM selection in SE tasks. The selection of an LLM for SE tasks should involve careful consideration rather than arbitrary choice. Key factors guiding this selection encompass the model’s proficiency in understanding the context of code, its ability to generate relevant content, responsiveness to fine-tuning, and demonstrated performance on SE-specific benchmarks (Xie et al., 2023; Li et al., 2023b, a). Given the stringent syntactical rules and functional requirements inherent to SE tasks, models capable of seamlessly integrating these complex aspects were typically favored.
@@
Task-specific fine-tuning. A notable trend is the customization of LLMs for precise SE tasks (Izadi et al., 2022; Li et al., 2023c; Zhang et al., 2022c). By fine-tuning models with datasets tailored to specific functions such as bug detection or code review, researchers were able to achieve marked performance improvements (Ciborowska and Damevski, 2023; Kou et al., 2023a).
@@
RQ1 - Summary (1) There are more than 50 different LLMs used for SE tasks in the papers we collected. Based on the underlying architecture or principles of different LLMs, we classified the summarized LLMs into three categories, i.e., encoder-only, encoder-decoder, and decoder-only LLMs. (2) We analyzed the trend of LLM usage for SE tasks. The most widely used LLMs are with decoder-only architectures. There are over 30 LLMs in the decoder-only category and 138 papers have researched the application of decoder-only LLMs to SE tasks.

### v6

@@ 2.3.2. Study Quality Assessment
A well-crafted quality assessment can help to prevent biases introduced by low-quality studies and can indicate to readers where caution about conclusions should be drawn (Yang et al., 2021). We formulated ten Quality Assessment Criteria (QAC), as shown in Table 4. These aim to assess the relevance, clarity, validity, and significance of included papers. We used a scoring system of -1, 0, 1 (irrelevant/unmet, partially relevant/met, relevant/fully met). The first three questions were designed for the remaining 382 papers in the fifth stage. If QAC1, QAC2, or QAC3 received a score of -1, there is no need to proceed with QAC4-QAC10, and the paper can be excluded directly. QAC4-QAC10 involved assessing the content of the papers using a scoring system of 0, 1, 2, 3 (poor, fair, good, excellent). Finally, we calculated the total score of QAC4-QAC10 for each paper. For published papers, the maximum score for QAC4-QAC10 should be 21 (3 $\times$× 7). We retained papers with a score of 16.8 (21 $\times$× 0.8) or above. For unpublished papers on arXiv, the score for QAC4 is always 0, and the maximum score for QAC5-QAC10 should be 18 (3 $\times$× 6). We retained papers with a score of 14.4 (18 $\times$× 0.8) or above. After this quality assessment, we obtained a final set of 382 papers.
@@ 2.4. Snowballing Search
Before conducting snowballing, a set of initial papers needs to be prepared. In this study, the initial paper list consists of the remaining 382 papers after the quality assessment. We performed forward and backward snowballing, which resulted in the collection of 3,964 and 9,610 papers, respectively. After initial deduplication, we were left with 5,152 papers. We then conducted the full study selection process on these 5,152 papers, including deduplicating them with the 382 papers from performing snowballing on the initial list. As a result, we obtained an additional 13 papers.
@@ 2.5. Data Extraction and Analysis
Figure 2. Overview of the selected 395 papers’ distribution.
@@ 4 The developed strategies and solutions
We finally obtained 395 relevant research papers after searching and snowballing. Fig. 2 presents an overview of the distribution of the included papers. As shown in Fig. 2 (a), 154 papers are published in peer-reviewed venues. ICSE is the most common of these venues, with a contribution of 41 papers. Other venues with noteworthy contributions include TSE, ESEC/FSE, and TOSEM, contributing 14, 12, and 11 papers respectively. Meanwhile, the remaining 241 papers are published on arXiv, an open-access platform that serves as a repository for scholarly articles. This finding is not surprising since much new LLM4SE research is rapidly emerging and thus many works are just completed and are likely in the peer review process. Despite the non-peer-reviewed nature of these papers, we have performed a rigorous quality assessment process on all collected papers, to ensure the quality of validity of our findings. This approach allows us to include all high-quality and relevant publications while maintaining high research standards.
Fig. 2 (b) shows the temporal distribution of the included papers. The number of publications has seen a rapidly growing trend since 2020. In 2020 and 2021, there are only 7 and 13 relevant papers, respectively. However, by 2022, the number of papers has increased dramatically to 56. What’s surprising is that, in 2023 alone, the number of published papers has already reached 273. And within just one month in 2024, 46 relevant papers are published. This rapid growth trend demonstrates that there is a growing research interest in the domain of LLM4SE.
In order to visualize the main content of our collection of papers, we generated a word cloud based on the abstracts of 395 papers as shown in Fig. 3. The most frequently occurring words include “code”, “LLM”, “language”, “model”, “large”, “task”, “software”,“generation”, “performance”, and “program”, clearly indicating the main themes explored in these papers. The terms “code” and “software” emphasize the core elements of software engineering, while “LLM”, “large”, “language” and “model” denote the use of large language models in a variety of tasks. The terms “generation”, “task”, and “program” emphasize the use of the LLM for automatic code generation and other SE tasks. In addition, “performance” reflects the evaluation and assessment of the effectiveness of LLM in SE applications. The word cloud provides further visual evidence that the literature we have collected is closely related to our research topic.
We then conducted data extraction during the full-text review. This extraction phase collected all relevant data that would facilitate a comprehensive and insightful response to the RQs outlined in Section 2.1. As depicted in Table 5, we extracted data including the classification of SE tasks, their corresponding activities, as well as the category, characteristics, and applicability of the LLMs. With this collected data, we systematically analyzed the relevant aspects of LLM4SE.
@@ 3.1. Large Language Models (LLMs)
Pre-trained language models (PLMs) have demonstrated impressive capabilities in solving various NLP tasks (Kojima et al., 2022; Shanahan, 2022; Wei et al., 2022b; Zhao et al., 2023d). Researchers have observed that scaling up the model sizes significantly enhances their capacity, leading to remarkable performance improvements when the parameter scale surpasses a certain threshold (Shanahan, 2022; Hoffmann et al., 2022; Taylor et al., 2022). The term “Large Language Model” (LLM) was introduced to distinguish language models based on their parameter size, specifically referring to large-sized PLMs (Zhao et al., 2023d). However, we note that the literature lacks a formal consensus on the minimum parameter scale for LLMs, as the model’s capacity is intertwined with both data size and total compute (Wang et al., 2023c). In this paper, we adopt the LLM scope division and taxonomy introduced by Pan et al.(Pan et al., 2023b) and categorize the mainstream LLMs investigated in this study into three groups according to their architectures: encoder-only, encoder-decoder, and decoder-only LLMs. This taxonomy and relevant models are shown in Fig. 4. We have included the LLMs used by each work and their parameter sizes (if declared in the paper) in our public repository: https://github.com/xinyi-hou/LLM4SE_SLR. Additionally, Table 6 summarizes the LLMs with different architectures suitable for different types of SE tasks.
@@
Table 6. Summary of LLMs with different architectures used in SE tasks.
Model Type Example of SE tasks
Encoder-only Understanding Code Understanding
Bug localization
Vulnerability detection
Encoder-Decoder Understanding and Generation Code summarization
Code translation
Program repair
Decoder-only Generation Code generation
Code completion
Test case generation
@@
Encoder-only LLMs. Encoder-only LLMs are a type of neural network architecture that utilizes only the encoder component of the model (Devlin et al., 2018). The encoder’s function is to process and encode the input sentence into a hidden representation, capturing the relationships between words and the overall context of the sentence. Notable instances of encoder-only LLMs include BERT (Devlin et al., 2018) and its variants (Feng et al., 2020; Guo et al., 2020; Liu et al., 2019; Lan et al., 2019). As an example, BERT’s structure, based on the Transformer’s encoder architecture, has been referenced in 50 our selected primary studies. Its distinctive bidirectional attention mechanism simultaneously considers the left and right context of each word during training. In the SE domain, other prominent models like CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2020), RoBERTa (Liu et al., 2019), and ALBERT (Lan et al., 2019) have been widely employed. Specialized models such as BERTOverflow (Tabassum et al., 2020) and CodeRetriever (Li et al., 2022b) have been specifically developed for SE applications. These models differ from BERT by leveraging program structure, introducing new pre-training tasks, or engaging new modalities, thereby improving the architecture’s application to code-related tasks. For example, CodeBERT integrates a token prediction scheme to comprehend code by predicting subsequent tokens, enhancing its understanding of programming languages for tasks like code completion and bug detection (Feng et al., 2020). GraphCodeBERT introduces edge-type prediction, recognizing relationships between code elements as a graph. This enables GraphCoderBERT to leverage code structure, improving its effectiveness in tasks like code summarization and program analysis (Guo et al., 2020). Encoder-only LLMs have shown efficacy in tasks requiring a nuanced understanding of the entire sentence or code snippet. Examples include code review, bug report understanding, and named entity recognition pertaining to code entities (Pudari and Ernst, 2023; Sghaier and Sahraoui, 2023; Yang et al., 2022c; Arakelyan et al., 2023; Li et al., 2023i; Mukherjee and Hellendoorn, 2023).
@@
Encoder-decoder LLMs. Encoder-decoder LLMs incorporate both encoder and decoder modules (Vaswani et al., 2017). The encoder ingests the input sentence and encodes it into a hidden space, effectively capturing the underlying structure and semantics. This hidden representation serves as an intermediary language, bridging the gap between diverse input and output formats. Conversely, the decoder utilizes this hidden space to generate the target output text, translating the abstract representation into concrete and contextually relevant expressions. Models such as PLBART (Ahmad et al., 2021), T5 (Raffel et al., 2020), and CodeT5 (Wang et al., 2021a) embodies this architecture. Further advancements are evident in CodeT5+ (Wang et al., 2023e), while AlphaCode (Li et al., 2022a) and CoTexT (Phan et al., 2021) showcase the architecture’s adaptability to various SE tasks. The encoder-decoder design offers flexible training strategies and is proficient in handling multifaceted tasks such as summarization, translation, and question-answering. Within the field of SE, this ability has been successfully applied to tasks like code summarization (Al-Kaswan et al., 2023; Gu et al., 2022; Mastropaolo et al., 2021b). The encoder module’s capacity to understand and represent both the structure and semantics of code is pivotal, allowing the decoder to translate this comprehension into concise, human-readable summaries.
@@
Decoder-only LLMs. Decoder-only LLMs exclusively utilize the decoder module to generate the target output text, following a distinct training paradigm that emphasizes sequential prediction (Radford et al., 2018). Unlike the encoder-decoder architecture, where the encoder processes input text, the decoder-only architecture begins with an initial state and predicts subsequent tokens, gradually building the output text. This approach relies heavily on the model’s ability to understand and anticipate language structure, syntax, and context. GPT-series models, such as GPT-1 (Radford et al., 2018), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), GPT-3.5 (OpenAI, 2022b), GPT-4 (OpenAI, 2023b), as well as their notable derivative, ChatGPT (OpenAI, 2022a) [^3] ChatGPT is a conversational agent built upon the GPT architecture, with GPT-3.5 and GPT-4 being specific versions of the architecture, each representing successive advancements., represent their major implementations. More specialized versions like CodeGPT (Lu et al., 2021), InstructGPT (Ouyang et al., 2022), Codex (Chen et al., 2021b), Copilot (GitHub, 2023) [^4] Copilot is an application built upon LLMs tailored for coding tasks. For convenience, all subsequent references in this paper to LLMs and their applications, such as ChatGPT and Copilot, will collectively be referred to as LLMs., and others have been fine-tuned for specific tasks in SE. Open-source models like GPT-J (Wang and Komatsuzaki, 2021), GPT-Neo (Black et al., 2021), GPT-NeoX (Black et al., 2022), LLaMA (Touvron et al., 2023a), and Vicuna (Chiang et al., 2023) also follow this architecture. Decoder-only LLMs are usually more suitable for various generation tasks, such as code generation and code completion. These models can generally perform downstream tasks from a few examples or simple instructions without adding prediction heads or fine-tuning, making them valuable tools in SE research. 2022 marked a surge in the development of decoder-only LLMs, a trend that gained further momentum in 2023, notably with the launch of commercial products by leading Internet companies. For example, Google launched Gemini (Google, 2023), Meta introduced LLaMA (Touvron et al., 2023a) and Llama 2 (Touvron et al., 2023b), and Anthropic unveiled Claude (Anthropic, 2023), etc. Contrary to LLMs such as GPT-4 and its derivative application, ChatGPT, released by OpenAI, which were promptly integrated into SE tasks, these new additions have not yet found widespread application within the SE field. Their potential remains largely unexplored, with opportunities for further assessment and utilization in specific tasks and challenges. The continued advancement of these models emphasizes the active exploration and innovation within decoder-only architectures.
@@ 3.2. Trend Analysis
As shown in Fig. 5, in the span from 2020 to 2024, the architecture of LLMs has witnessed notable shifts in preference and application within SE tasks. The specific choices between decoder-only, encoder-decoder, and encoder-only structures have shaped the direction of research and solutions in the SE domain (Wong et al., 2023). This analysis explores trends in the adoption of these architectures over the years, reflecting the evolving dynamics of LLM for SE tasks.
@@
Evolution of LLM architectures in 2021. The year 2020 saw research papers predominantly concentrating on encoder-only LLMs for SE tasks, evidenced by a total of eight papers. Decoder-only LLMs or encoder-decoder LLMs were scarcely featured in that year’s research. A marked change occurred in 2021. Out of 19 papers in 2021, nine were dedicated to decoder-only LLMs, constituting 47.37% of the research. Additionally, two papers, or 10.53%, focused on encoder-decoder LLMs. Encoder-only LLMs witnessed a slight decline, representing 42.1% of the field with eight papers. This rapid transition can be linked to the generative capability of decoder-only LLMs. Researchers (Laskar et al., 2023; Sadik et al., 2023; Sridhara et al., 2023) found that these models, e.g., GPT series, requiring minimal fine-tuning, could produce not only syntactically correct but also functionally relevant code snippets. Their proficiency in grasping the context of code quickly made them a preferred choice.
@@
Diversity of LLM architectures in 2022. 2022 experienced a significant increase in diversity, with more varied LLM architectures finding representation. Out of a total of 142 papers, 73 were centered around decoder-only LLMs, comprising 51.41% of the studies. Encoder-decoder LLMs made their presence known in 17 papers, accounting for 11.97%. Meanwhile, encoder-only LLMs led the field slightly with 52 papers, capturing 36.62% of the research interest. This diverse distribution suggests an exploration phase where researchers were actively assessing and leveraging different architectures to suit varied needs and challenges. The near-equal interest across different architectures underscores the field’s richness, indicating that no single approach had become the definitive choice.
@@
Dominance of the decoder-only architecture in 2023. 2023 signaled a strong shift towards decoder-only LLMs. An impressive 432 instances of utilizing decoder-only LLMs were recorded across 195 unique papers, reflecting that a single paper might employ multiple such models. These papers focusing on decoder-only LLMs constituted a significant 70.7% of the total research this year. In comparison, encoder-decoder LLMs were the subject of 85 papers, contributing 13.91%, while encoder-only LLMs appeared to stabilize, with 94 papers, representing 15.39% of the 2023 research landscape. This trend signifies a shift in focus and resources toward exploring and harnessing the decoder-only architecture as the primary approach in many current and future LLM4SE research and applications.
Exploration of the LLM architecture in 2024. The initial trends in January 2024 showcase the ongoing evolution of LLM architectures. Among the 120 papers examined, decoder-only LLMs continued to maintain a prominent position, with 77 papers dedicated to this architecture, constituting 64.17% of the research. Encoder-decoder LLMs appeared in 24 papers, representing 20% of the total, while encoder-only LLMs were featured in 19 papers, making up 15.83%. Although there is a slight decrease in the dominance of decoder-only architectures compared to the previous year, they still hold a central role. The persistent exploration of encoder-decoder and encoder-only architectures suggests an enduring interest in diverse configurations within the SE research community.
@@
Criteria for LLM selection in SE tasks. The selection of an LLM for SE tasks should involve careful consideration rather than arbitrary choice. Key factors guiding this selection encompass the model’s proficiency in understanding the context of code, its ability to generate relevant content, responsiveness to fine-tuning, and demonstrated performance on SE-specific benchmarks (Xie et al., 2023a; Li et al., 2023c, b). Given the stringent syntactical rules and functional requirements inherent to SE tasks, models capable of seamlessly integrating these complex aspects were typically favored.
@@
Task-specific fine-tuning. A notable trend is the customization of LLMs for precise SE tasks (Izadi et al., 2022; Li et al., 2023i; Zhang et al., 2022c). By fine-tuning models with datasets tailored to specific functions such as bug detection or code review, researchers were able to achieve marked performance improvements (Ciborowska and Damevski, 2023; Kou et al., 2023a).
@@
RQ1 - Summary
(1) There are more than 70 different LLMs used for SE tasks in our selected primary studies. Based on the underlying architecture or principles of different LLMs, we classified the summarized LLMs into three categories, i.e., decoder-only, encoder-decoder, and encoder-only LLMs. (2) We observed that each LLM architecture serves a specific purpose in SE tasks, with encoder-only LLMs focusing on comprehensive understanding, encoder-decoder LLMs used for tasks requiring understanding of input information followed by content generation, and decoder-only LLMs being more suitable for generation tasks. (3) We analyzed the trend of LLM usage for SE tasks. The most widely used LLMs are with decoder-only architectures. There are over 45 LLMs in the decoder-only category and 195 papers have researched the application of decoder-only LLMs to SE tasks.